OpenStudy (anonymous):

A 7.6m bean is supported by two posts. Post A is on the left end of the beam. Post B is 2m to the right from post A. Them Beam weighs 1200 N. What forces must be exerted at A and B to maintain equilibrium? In what direction must each force act?

OpenStudy (anonymous):

First, let's draw a FBD. This is always FIRST and foremost. where w is the weight distribution of the beam, R is the resultant force of the distributed force, and x is the location from the left side of R. Take the reaction forces at A and B to be upwards. Second, we need to find the location of the reaction force. This is a simple case. The location of a reaction force coresponding to a square distributed force is half the length of the distribution, in this case, it is distributed over the length of the beam. \[x = {l \over 2} = {7.6 \over 2}\]The magnitude of R is simply the weight. \[R = 1200 ~\rm N\] Finally, let's sum forces and torques. \[\sum F_Y = 0 \rightarrow F_A + F_B - R = 0\]\[\sum \tau_A = 0 \rightarrow F_B \cdot 2 - R \cdot x = 0\]We can solve the torque equation for \(F_B\) and then solve the force equation for \(F_A\). Note that \(F_B\) will be directed upwards (positive) and \(F_A\) will be downwards (negative).